geerky42 (geerky42):

If a/6 has a remainder of 5 and b/6 has a remainder of 4, then (a + b)/6 has a remainder of _________.

OpenStudy (anonymous):

\[\frac{a+b}{6}=\frac{a}{6}+\frac{b}{6}\]

geerky42 (geerky42):

And? I don't know rocket science.

OpenStudy (shamim):

result will b 3

OpenStudy (shamim):

a=11 b=10

OpenStudy (shamim):

a/6 has remainders 5 so the value of a will b 11

OpenStudy (anonymous):

How did you get 11?

OpenStudy (shamim):

b/6 has remainders 4 so the value of b will b 10

OpenStudy (shamim):

6+5=11

OpenStudy (anonymous):

Oh god,

OpenStudy (anonymous):

I have no idea why you would even add them.

OpenStudy (shamim):

so (a+b)/6 =(11+10)/6 =21/6

OpenStudy (shamim):

it has remainder 3

OpenStudy (anonymous):

Ouch, Please don't write down three.

OpenStudy (anonymous):

jiji it's not saying there is a remainder of 5 the equation would look lik this \[\frac{a}{6}+\frac{5}{6}+\frac{b}{6}+\frac{4}{6}\]

OpenStudy (anonymous):

It IS saying there is a remainder of 5.

OpenStudy (anonymous):

since when is a remaind a whole number ?

OpenStudy (anonymous):

for example divide a polynomial by a smalller polynomial f(x). Do you just say the whole number as the remainder or do you say \[\frac{a}{f(x)}\]

OpenStudy (anonymous):

I think I'll wait for the answer.

OpenStudy (anonymous):

if the remainder was just 5, there wouldn't be a remainder: it would divide evenly

OpenStudy (anonymous):

Mhmm true.

OpenStudy (anonymous):

I gave the biggest hint and i actually showed what it would be

OpenStudy (anonymous):

a+5/6 then?

OpenStudy (anonymous):

just never gave the exact

OpenStudy (anonymous):

yes for the first... and then for the second

OpenStudy (anonymous):

We'll let the guy who wrote this question do it ;)

OpenStudy (anonymous):

so what you end up with is \[\frac{9}{6}=1+\frac{3}{6}\]

OpenStudy (shamim):

sorry

OpenStudy (anonymous):

now what you get is \[c_1+c_2+1+\frac{3}{6}\] where \[c_1,c_2 \] are whole \[(c_1+c_2+1)+\frac{3}{6}\]